OpenStudy (anonymous):

Find the equation of the line that passes through (2, 9) with a slope of 4. Write your answer in Slope-Intercept Form.

OpenStudy (anonymous):

y=mx+b 9=4(2)+b 9=8+b 1=b => y=4x+1